# Enhancing Memory Efficiency in Large Language Model Training Through Chronos-aware Pipeline Parallelism

### v1

@@ Abstract.
Larger model sizes and longer sequence lengths have empowered the Large Language Model (LLM) to achieve outstanding performance across various domains. However, this progress brings significant storage capacity challenges for LLM pretraining. High Bandwidth Memory (HBM) is expensive and requires more advanced packaging technologies for capacity expansion, creating an urgent need for memory-efficient scheduling strategies. Yet, prior pipeline parallelism schedules have primarily focused on reducing bubble overhead, often neglecting memory efficiency and lacking compatibility with other memory-efficient strategies. Consequently, these methods struggle to meet the storage demands of storage capacity for next-generation LLM.
This work presents ChronosPipe, a Chronos-aware pipeline parallelism for memory-efficient LLM pretraining. The core insight of ChronosPipe is to treat HBM as a fast but small ’cache,’ optimizing and exploiting temporal locality within LLM pretraining to enhance HBM utilization. ChronosPipe introduces a pipeline scheduling strategy, Chronos-Pipe, to reduce the extrinsic overhead that disrupts the temporal locality of activations. Additionally, it leverages Chronos-Recomp and Chronos-Offload to efficiently harness the intrinsic temporal locality of activations and weights in Deep Neural Networks. Experiment results show that ChronosPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. INTRODUCTION
Contemporary GPU training systems utilize high-bandwidth memory (HBM) for runtime variable storage, but it now faces significant challenges. First, HBM integration requires advanced packaging techniques such as through-silicon via (TSV) (TSV, ) and microbump (microbump, ), resulting in substantial manufacturing costs. For instance, HBM accounts for 50-60% of the total costs in NVIDIA’s H100 SXM5 module (H100, ). Moreover, next-generation HBM poses considerable problems. Increasing the count of stacking DRAM die within the restricted packaging height (720 um) necessitates hybrid bonding (HybridBonding, ), whose thermal management and yield optimization (HybridBonding_Problem, ) are extremely challenging.
@@
Figure 1. (a) limited HBM capacity is difficult to meet the demands of next-generation LLM. (b) existing pipeline parallelism ignores temporal locality
Moreover, current LLM training predominantly employs hybrid parallelism, yet existing PP strategies lack optimal compatibility with other memory-efficient scheduling approaches. The ZeRO series (ZeRO, ) represents state-of-the-art data parallelism (DP) techniques, distributing model states across DP machines. In hybrid DP and PP configurations, distributing optimizer states (ZeRO-1) necessitates inter-DP rank communication only at mini-batch granularity. However, further distributing weight gradients (ZeRO-2) and weights (ZeRO-3) requires more frequent communication at micro-batch granularity, imposing higher bandwidth demands between DP ranks. Consequently, hybrid DP-PP implementations (DeepSpeed, ) typically utilize only ZeRO-1, leaving substantial potential for reducing weight and gradient storage. When designing PP strategies that are more compatible with ZeRO-2/3, such as Breadth First PP (BreadthFirstPP, ) and ZeROPP (ZeroPP, ), these approaches reduce DP bandwidth requirements at the cost of a significant increase in activation memory, leaving no perfect solution.
@@
Unlike previous work, we notice the issues mentioned above and introduce the concept of temporal locality to pipeline parallelism for the first time. We propose ChronosPipe, which enables the training of models 2.4 times larger than 1F1B (DAPPLE, ) on the same hardware with comparable throughput while maintaining compatibility with various memory-efficient scheduling strategies. The ChronosPipe solution comprises three components: Chronos-Pipe, Chronos-Recomp, and Chronos-Offload, which are described below.
Chronos-Pipe minimizes peak activation storage by reducing micro-batch execution time, thus expediting the backward pass that consumes activations. This, in effect, optimizes activation temporal locality. Chronos-Recomp exploits the poorer temporal locality of shallower layer activations and discards them from HBM by selectively recomputing to achieve higher efficiency. Chronos-Offload takes advantage of poorer temporal locality of deeper layer weight and discards them from HBM by offloading the process of optimizer update to CPU. Since this process overlaps with PP’s warmup and cooldown phase, it lowers demand on offload bandwidth and the processing speed of CPU.
@@
  • We found that temporal locality in DNN is the primary cause of imbalanced activation storage in PP. This imbalance limits the size of trained models on current devices.
  • For activation memory savings, we introduce temporal locality into the PP schedule (Chronos-Pipe) and recomputation (Chronos-Recomp). Chronos-Pipe eliminates unnecessary intervals, optimizing the temporal locality of activation (Section 4.1). Chronos-Recomp discards activation with poor temporal locality from HBM by selectively recomputing shallow layers to achieve higher efficiency (Section 4.2).
  • For Model State storage capacity, we show that a ZeRO-2-compatible PP could built based on Chronos-Pipe (Section 4.3). Additionally, we designed the offload strategy based on temporal locality (Chronos-Offload), which can discard model states with poor temporal locality from HBM by offloading optimizer updates of deeper layers to the CPU (Section 5.1).
  • End-to-end evaluation on ChronosPipe is carried out on a cluster comprising up to 64 accelerators. Experiments show that Chronos-Pipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
@@ 2. BACKGROUND AND MOTIVATION
In this section, we will introduce pipeline parallelism (PP) (Section 2.1), typical recomputation strategies (Section 2.2), and offloading strategies (Section 2.3). We will then discuss the motivation for leveraging temporal locality (Section 2.4).
@@
$\bullet$ $P$:# of pipeline stages
$\bullet$ $m$:# of micro-batches executed within a training iteration
$\bullet$ $T_{fwd}$: execution time of forward pass in one micro-batch
$\bullet$ $T_{bwd}$: execution time of backward pass in one micro-batch
@@ 2.1. Pipeline parallelism
Figure 2. (a) Interleave-1F1B schedule. (b) ZeRO-compatible pipeline schedule
@@
Pipeline Parallelism (PP) divides the model into multi-layer blocks mapped across different PP stages. These stages work together to train all samples within a mini-batch, enabling mini-batch Stochastic Gradient Descent. During the forward pass, samples in each mini-batch begin computation at the first stage, with subsequent blocks executing their computations in a pipelined manner until the final forward computation completes at the last stage. In the backward pass, gradients for each mini-batch flow in the reverse direction. As a result, PP only requires a single tensor to be sent and received per block, significantly reducing communication bandwidth demands. This approach is often deployed across nodes and is a key component in hybrid parallelism.
@@
Pipeline parallelism (PP) at the mini-batch level suffers from significant pipeline bubbles, prompting the common practice of splitting mini-batches into smaller micro-batches. By dividing the workload, the first stage, after processing the initial micro-batch, can immediately start on subsequent micro-batches rather than idling, thus reducing pipeline bubbles (GPipe, ). When the number of micro-batches increases, the classic 1F1B (one-forward-one-backward) scheduling strategy (DAPPLE, ) is often used, as shown in Fig.  1(b). During the steady phase, a forward computation block is initiated only after a backward block finishes, avoiding the simultaneous launch of multiple forward passes in the warm-up phase, which would otherwise increase peak memory usage. However, this approach creates an imbalance in activation memory usage across PP stages. In a single micro-batch, the first stage starts forward computation early but is the last to complete backward passes and release memory. As a result, the first stage reaches peak activation memory of $m_{a}$, while the last stage’s activation memory usage is only $\frac{m_{a}}{p}$. Additionally, Fig.  1(b) illustrates that tasks with dependencies between adjacent PP stages execute sequentially, leading to non-overlapping point-to-point (P2P) communication. Some scheduling methods address this by inserting independent task blocks from other micro-batches between dependent blocks, enabling overlapping P2P communication. This delay caused by dependency constraints is called the ”interval” delay in this paper.
@@
Blocks can be broken down into smaller tasks to reduce pipeline bubbles further, though this may increase memory requirements. The Interleaved-1F1B approach (Interleave-1F1B, ) divides blocks into more chunks and schedules multiple chunks launches during the warmup phase, as shown in Fig. 2(a). When the workload is divided into v chunks (assigning layer 1, $p+1$, …, $(v-1)p+1$ to stage 0; layer 2, $p+2$, …, $(v-1)p+2$ to stage 1, and so forth), the bubble size is reduced to $\frac{1}{v}$ of that in the traditional 1F1B, but peak activation memory increases to $m_{a}(1+\frac{p-1}{pv})$. The Zero Bubble approach (ZeroBubble, ) goes a step further by splitting the backward pass (BP) into two phases: activation gradient computation (BPA) and weight gradient computation (BPW). This arrangement theoretically enables a bubble-free pipeline. However, the activation gradients generated by BPA are required to be retained until BPW is completed, which raises memory demands. Additionally, splitting the BP disrupts the original reuse of activation gradients for all-gather operations between BPA and BPW, and reduces the overlap between computation and communication.
Designing a ZeRO-compatible pipeline schedule is challenging due to the trade-off between DP communication bandwidth and activation memory storage. Communication between DP ranks at the micro-batch level is required for a schedule to be compatible with ZeRO-2 or ZeRO-3. To overlap this communication, existing methods, such as Breadth-First PP (BreadthFirstPP, ) and ZeROPP (ZeroPP, ), use grouped execution of micro-batches, as shown in Fig. 2(b). This approach processes multiple micro-batches simultaneously, allowing communications of weights and weight gradients can be overlapped with computations, but it also causes a substantial increase in activation memory usage. Since weights and weight gradients are stored in 16-bit and 32-bit precision, respectively, achieving ZeRO-3 PP requires at least a 1.5x increase in DP communication bandwidth or an increase in the number of micro-batches per group compared to ZeRO-2 PP. Consequently, some implementations choose to use only ZeRO-2 PP (LLAMA3, ).
2.2. Recomputation
@@
Recomputation allows only a subset of activations generated in the forward pass (called checkpoints) to be retained, with all other activations regenerated during the backward pass, trading off memory usage for additional computation. Previous work (ChainTrainingMem, ) has explored optimal checkpoint placement, given a fixed number of checkpoints, to minimize activation memory, achieving memory savings at sub-linear cost for chain structure of DNN. Additionally, the benefits of recomputation vary across different operators. Recent studies suggest recomputing only the attention part in transformers (SeqParallel, ) or recomputing some non-linear operators (MegatronKwai, ) to reduce memory usage with minimal computational overhead effectively. So far, it is still expensive to recompute the projection operator in LLM, which dominates the remaining activation memory usage. Moreover,  (DTR, ) further introduced temporal locality-aware recomputation principle and demonstrated its efficacy in data parallelism. Although pipeline stages inherently exhibit temporal locality variations in PP, coarse-grained PP fails to leverage intra-block temporal locality variance for recomputation optimization within the simplistic structure of LLMs. Existing methods thus fail in single-chunk PP. Moreover, Direct application of this principle to multi-chunk PP (e.g., interleaved 1F1B scheduling) risks inducing steady-phase bubbles, resulting in significant throughput degradation.Therefore,  (DTR, ) fail to solve imbalanced memory in PP.
2.3. Offloading Strategy
@@
Offloading techniques allow tensors to be moved from memory-constrained GPUs to other GPUs or CPUs, trading memory usage for data movement. In BPipe (Bpipe, ), GPU offloading is used to address activation memory imbalances in pipeline parallelism. However, this approach requires high-bandwidth intra-node communication links like NVLink rather than slower inter-node links like Ethernet or InfiniBand. CPU offloading utilizes PCIe to offload activation or model states. Theoretically, CPU offloading for activation memory reduction is well-suited for long-sequence training (MegatronKwai, ). However, limited PCIe bandwidth may leave some activation in HBM in practice. For model states, previous work, such as ZeRO-Offload (ZeRO-Offload, ), offloads the process of optimizer-step to the CPU, storing both optimizer states and master weights in CPU-side DRAM. To minimize GPU idle time caused by latency in the CPU’s optimizer step, this approach imposes significant requirements on both offload bandwidth and CPU computational capacity. Overall, Recent studies still have high bandwidth requirements for the Offloading Strategy.
2.4. Motivation: Temporal-Locality Matters
@@
Higher Peak Activation Storage Requirements: The PP stage responsible for shallow-layer activations releases them last, which increases the activation storage requirements. Additionally, intervals can exist between blocks with dependencies, and these intervals can accumulate across multiple blocks, further worsening the temporal locality for shallow layers. This is the reason that the interleaved 1F1B schedule results in a higher peak storage demand.
@@
Skewed Distribution of Peak Activation Storage: As shown in Fig.  2(a), the peak activation storage in the interleaved 1F1B schedule tends to occur at Stage 0. When the chunk size is set to 2, the ratio of peak activation storage between shallow and deep layers is $(2p-1):p$, indicating an obvious bias.
@@
Sufficient Time for Process of optimizer-step: Beyond the natural advantage of chain-structured networks providing additional time for optimizer update on deep-layer weights, increasing the PP will provide more sufficient time for optimizer-step. Large PP does not impact the execution latency of each micro-batch while the amount of model state that needs updating per stage is reduced, thereby lowering the communication and processing demands for weight updates.
3. OVERVIEW OF ChronosPipe

### v2

@@ Abstract.
Scaling LLMs through increased parameters and sequence lengths imposes significant storage pressure on LLM training. While expensive HBM requires advanced packaging for capacity scaling, existing pipeline parallelism strategies prioritize computational bubble reduction over memory efficiency. Consequently, existing methods fail to address the increasing storage requirements for ever-growing LLMs.
This work presents TPipe, a temporal-locality-aware pipeline parallelism for memory-efficient LLM training. The core insight of TPipe is to treat HBM as a fast but small ’cache,’ optimizing and exploiting temporal locality within LLM. TPipe solution comprises three components: T-Pipe, T-Recomp, and T-Offload. T-Pipe, a pipeline scheduling strategy, is used to reduce the extrinsic overhead that disrupts the temporal locality of activations. T-Recomp and T-Offload are utilized to efficiently harness the intrinsic temporal locality of the activations and weights in DNN. Experiment results show that TPipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. INTRODUCTION
Contemporary GPU training systems utilize high-bandwidth memory (HBM) for runtime variable storage, but it now faces significant challenges. First, HBM integration requires advanced packaging techniques such as through-silicon via (TSV) (TSV, ) and microbump (microbump, ), resulting in substantial manufacturing costs. For instance, HBM accounts for 50-60% of the total costs in NVIDIA’s H100 SXM5 module (H100, ). Moreover, increasing the count of stacking DRAM die within the restricted packaging height (720 um) necessitates hybrid bonding (HybridBonding, ), whose yield optimization (HybridBonding_Problem, ) is extremely challenging.
@@
Figure 1. (a) limited HBM capacity is difficult to meet the demands of next-generation LLM. (b) existing pipeline parallelism (such as 1F1B) ignores temporal locality.
@@
Unlike previous work, we notice the issues mentioned above and introduce the concept of temporal locality to pipeline parallelism for the first time. We propose TPipe, which enables the training of models 2.4 times larger than 1F1B (DAPPLE, ) on the same hardware with comparable throughput while maintaining compatibility with various memory-efficient scheduling strategies. The TPipe solution comprises three components: T-Pipe, T-Recomp, and T-Offload.
T-Pipe minimizes peak activation storage by reducing micro-batch execution time, thus expediting the backward pass that consumes activations. This, in effect, optimizes activation temporal locality. T-Recomp exploits the poorer temporal locality of shallower layer activations and discards them from HBM by selectively recomputing to achieve higher efficiency. T-Offload takes advantage of poorer temporal locality of deeper layer weight by offloading optimizer update to CPU during bubbles created in T-Pipe. Since this process overlaps with warmup/cooldown phase, it lowers demand on offload bandwidth and CPUs. These strategies co-design based on temporal locality variations in pipeline parallelism, which remain underexplored in existing methods.
@@
  • We found that temporal locality in DNN is the primary cause of imbalanced activation storage in PP, constraining trainable model sizes on current devices.
  • For activation memory savings, we introduce temporal locality into the PP schedule (T-Pipe) and recomputation (T-Recomp). T-Pipe eliminates unnecessary intervals, optimizing the temporal locality of activation (Section 4.1). T-Recomp discards activation with poor temporal locality from HBM by selectively recomputing to achieve higher efficiency (Section 4.2).
  • For model state storage, we designed the offload strategy based on temporal locality (T-Offload), which can discard model states with poor temporal locality from HBM by offloading optimizer updates of deeper layers to CPU during bubbles created in T-Pipe (Section 5).
  • End-to-end evaluation on TPipe is carried out on a cluster comprising up to 64 GPUs. Experiments show that T-Pipe can expand the trainable model size by 2.4x while maintaining comparable throughput, achieving 1.5x better than the 1F1B strategy combined with recomputation.
@@ 2. BACKGROUND AND MOTIVATION
In this section, we will introduce pipeline parallelism (PP) (Section 2.1), typical recomputation strategies and offloading strategies (Section 2.2). We will then discuss the motivation for leveraging temporal locality (Section 2.3) and review how prior work has addressed memory imbalance challenges in PP (Section 2.4).
@@
$\bullet$ PP:# of pipeline stages
$\bullet$ mm:# of micro-batches within a training iteration
$\bullet$ $T_{fwd}/T_{bwd}$: execution time of forward/backward pass in one micro-batch
@@ 2.1. Pipeline parallelism
Figure 2. (a) Interleave-1F1B (b) and its memory usage
@@
Pipeline Parallelism (PP) divides the model into multi-layer blocks mapped across different PP stages. These stages work together to train all samples within a mini-batch, enabling mini-batch Stochastic Gradient Descent. Forward computations in each mini-batch propagate sequentially through stages until completion, while gradients flow inversely during backward passes. As a result, only one layer of intermediate activation/gradient needs to be sent and received per block, significantly reducing communication bandwidth demands.
@@
Mini-batch-level PP suffers from significant pipeline bubbles, commonly addressed by splitting mini-batches into micro-batches. Dividing workloads lets the first stage proceed from initial to subsequent micro-batches without idling, thus reducing pipeline bubbles (GPipe, ). When the number of micro-batches increases, the classic 1F1B strategy (DAPPLE, ) is often used, as shown in Fig.  1(b). During the steady phase, each forward computation block follows a backward block completion, preventing concurrent forward passes during warm-up that elevate peak memory usage. However, this approach creates an imbalance in activation memory usage across PP stages. In a single micro-batch, the first stage initiates forward computation first but finalizes backward passes and release memory last. As a result, the first stage reaches peak activation memory of $m_{a}$, while the last stage’s activation memory usage is only $\frac{m_{a}}{p}$. Fig.  1(b) illustrates that tasks with dependencies between adjacent PP stages execute sequentially, leading to non-overlapping point-to-point (P2P) communication. Some methods mitigate this by inserting independent tasks from other micro-batches between dependent blocks to enable communication overlap. This delay caused by dependency is called the ”interval” in this paper.
@@
Blocks can be broken down into smaller tasks to reduce pipeline bubbles further, though this may increase memory requirements. Interleaved-1F1B  (Interleave-1F1B, ) divides blocks into more chunks and schedules multiple chunks launches during the warmup phase, as shown in Fig. 2(a). When the workload is divided into v chunks (assigning layer 1, $p+1$, …, (v1)p+1(v-1)p+1 to stage 0; layer 2, $p+2$, …, (v1)p+2(v-1)p+2 to stage 1, and so forth), the bubble size is reduced to $\frac{1}{v}$ of that in the traditional 1F1B, but peak activation memory increases to $m_{a}(1+\frac{p-1}{pv})$ (Fig. 2(b)). Zero Bubble (ZeroBubble, ) splits backward propagation (BP) into activation gradient (BPA) and weight gradient (BPW), enabling a nearly ’bubble-free’ pipeline. BPA-generated gradients must persist until BPW completion, which raises memory demands. Additionally, splitting the BP disrupts reuse of activation gradients for all-gather operations between BPA and BPW, and reduces the overlap between computation and communication.
2.2. Recomputation and Offloading Strategy
@@
Recomputation allows only a subset of activations generated in the forward pass (called checkpoints) to be retained, with all other activations regenerated during the backward pass, trading off memory usage for additional computation. Previous work (ChainTrainingMem, ) introduced recomputation into DNNs, proving that optimal scheduling enables memory savings at sublinear cost for chain-structured DNN. Additionally, the benefits of recomputation vary across different operators. Recent studies suggest recomputing only the attention part in transformers (SeqParallel, ) or recomputing some non-linear operators (MegatronKwai, ) to reduce memory usage with minimal computational overhead effectively. So far, it is still expensive to recompute the projection operator in LLM, which dominates the remaining activation memory usage. Moreover,  (DTR, ) further introduced temporal locality-aware recomputation principle and demonstrated its efficacy in data parallelism. However, temporal locality-aware recomputation in pipeline parallelism remains underexplored, necessitating systematic solutions for inter-stage dependency management and programming complexity mitigation.
@@
Offloading techniques allow tensors to be moved from memory-constrained GPUs to other GPUs or CPUs, trading memory usage for data movement. In BPipe (Bpipe, ), GPU offloading is used to address activation memory imbalances in pipeline parallelism. However, this approach requires high-bandwidth intra-node communication links like NVLink. CPU offloading utilizes PCIe to offload activation or model states. Theoretically, CPU offloading for activation memory reduction is well-suited for long-sequence training (MegatronKwai, ). However, completely offloading all the activations can lead to significant performance degradation due to limited PCIe bandwidth. For model states, previous work, such as ZeRO-Offload (ZeRO-Offload, ), offloads optimizer-step to CPUs. To minimize GPU idle time caused by latency in the CPU’s optimizer step, this approach imposes significant requirements on both offload bandwidth and CPU computational capacity. Overall, Recent studies still have high bandwidth requirements for the Offloading Strategy.
2.3. Motivation 1: Temporal-Locality Matters
@@
Higher Peak Activation Storage Requirements: The PP stage responsible for shallow-layer activations releases them last, which increases the activation storage requirements. Additionally, intervals can exist between blocks with dependencies, and these intervals can accumulate across multiple blocks, further worsening the temporal locality.
@@
Skewed Distribution of Peak Activation Storage: As shown in Fig.  2, the peak activation storage in the interleaved 1F1B tends to occur at Stage 0. When the chunk size is set to 2, the ratio of peak activation storage between shallow and deep layers is (2p1):p(2p-1):p, indicating an obvious bias.
@@
Sufficient Time for optimizer-step: Beyond the natural advantage of chain-structured networks providing additional time for optimizer update on deep-layer weights, increasing the PP will provide more sufficient time for optimizer-step. Larger pipeline parallelism does not impact execution latency of each micro-batch while reducing amount of per-stage model state updates. Additionally, warm-up/cool-down phase bubbles can be strategically utilized to reduce communication and processing overhead during weight updates.
2.4. Motivation 2: Memory-efficient PP Solutions
To address the memory imbalance problem in pipeline parallelism (PP), prior work primarily follow two directions. The first way employs stage-aware strategies like AdaPipe (AdaPipe, ), which dynamically allocates varying computational workloads and recomputation ratios across pipeline stages to achieve memory balance. However, this adaptive allocation introduces significant compilation and programming complexity as it must continuously adjust to different model architectures and hardware configurations.
The second approach adopts locality-balanced solutions exemplified by V-Half/V-Min  (CtrlMemPP, ), which utilize a V-shaped scheduling pattern to assign tasks with the best temporal locality (deepest layer) and worst temporal locality (shallowest layer) within the same pipeline stage, thus statically balancing memory across stages while maintaining uniform computational loads.
Despite their merits, V-Half/V-Min exhibit critical limitations. First, while achieving stage-level memory balance through V-shaped scheduling, these methods still retain tensors with poor temporal locality, failing to reach the upper bound of memory optimization via temporal locality. Second, their memory-efficient V-shaped design requires partitioning backpropagation (BP) into separate BPA and BPW phases, preventing overlap between tensor parallelism’s collective communications and computation. This inefficiency becomes particularly severe in long-sequence training, where it introduces persistent pipeline bubbles and degrades computational efficiency.
Unlike previous work, we propose TPipe, a rigorous temporal-locality-aware design to systematically address these limitations, as detailed in the following sections.